OpenStudy (willgrant10):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)

Angle:

Hello there! :) Point-slope form of a line is in the format of ... y - k = m(x - h) or you might have also learned it as \(y - y_1 = m(x - x_1)\) both are the same :)

Angle:

the h and k come from a given point of the format (h, k) same with the \(x_1\) and \(y_1\) which come as \((x_1 , y_1)\) in this case, the given points are (-2, -5) and (2, 3) either one of these points would work for our (h, k) or \((x_1, y_1)\)

Angle:

m in the equation is the slope > same as it was in the slope-intercept form (y=mx+b) that I'm assuming you have already learned ^_^"

Angle:

Now that we have covered what the point-slope form is and what each part of it is, we can proceed to try to fill in the form :)

Angle:

the first step is to find the slope (m) \(\large {slope} = \huge {\frac{y_2 - y_1}{x_2 - x_1}}\) when given the points \((x_1,y_1)\) and \((x_2,y_2)\) so when we have (-2, -5) and (2, 3) we can easily plug it into this slope function and get our slope :)

Angle:

knowing all this, you can try plugging either of the points given in the question and the slope you find and get your answer ^_^ if you still need help, feel free to tag me with @Angle ! Good Luck :)